OpenStudy (elisaneedshelp):

What is the solution set for the equation x+2= √4x+13 ? Pleasee Medal and Fan

OpenStudy (elisaneedshelp):

\[x+2=\sqrt{4x+13}\]

OpenStudy (elisaneedshelp):

a. (-3,2) b. (-3,3) c. (-2,2) d. (-2,3)

Nnesha (nnesha):

how would you cancel square root ???

OpenStudy (elisaneedshelp):

I believe the 4x would turn into a 2x but then i wouldnt know what to do with the 13

Nnesha (nnesha):

x have an exponent which is one so you can't take square root of that to cancel out square root you should take square both side

OpenStudy (elisaneedshelp):

can you show me? :) please sorry

Nnesha (nnesha):

 that's how you should take square both side remember exponent rules!! you can change square root to exponent \[\large\rm \sqrt{x} = x^\frac{ 1 }{ 2 }\] so when you multiply that by 2 both 2/2 can cancel out